# Magnetorheological Characterization of Blood Analogues Seeded with Paramagnetic Particles

### v1

@@ 3.3 Numerical analysis
Having the critical shear rates, we could define a strain rate range to test in oscillatory flow for either chain length. We aimed to test one strain rate for stable chains ($\dot{\gamma}_{0}<\dot{\gamma}_{\mathrm{sub.}}$), two within the intermediate range ($\dot{\gamma}_{\mathrm{sub.}}<\dot{\gamma}_{0}<\dot{\gamma}_{\mathrm{brk.}}$), and two past the breakdown threshold ($\dot{\gamma}_{0}>\dot{\gamma}_{\mathrm{brk.}}$ and $\dot{\gamma}_{0}\gg\dot{\gamma}_{\mathrm{brk.}}$). As such, for the chains with $l_{\mathrm{chain}}=4$ and 10, we defined, respectively: $\dot{\gamma}_{0}\in[0.1,0.3,0.7,1,5]$ s-1 and $\dot{\gamma}_{0}\in[0.01,0.1,0.7,1,5]$ s-1. Ten strain values were tested between $0.1\leq\gamma_{0}\leq 1$, running for 3 cycles ($t=3\times 2\pi/\omega$, sufficient to achieve a steady cycle) and the flow was driven by a top wall velocity $U=\dot{\gamma}_{0}\,h_{\mathrm{c}}\,\cos(\omega\,t)$, with $\omega=\dot{\gamma}_{0}/\gamma_{0}$.
@@
Having the particle positions, we want to probe for chain subdivision. In steady shear we could simply analyse the interparticle distance as post chain-break the particles continuously diverge, but in oscillatory shear this method alone is insufficient due to the time-dependent nature of the flow and because the minimum distance for chain break is somewhat arbitrary. Here we choose to consider subdivision if, within the final simulated cycle, the distance between the central particle-pair exceeds 10% of their aggregated distance, $r/d_{\mathrm{p}}>1.1$, and, additionally, if the critical angle is surpassed $\phi>33.5^{\circ}$. In Figure 10 are shown the passing conditions for chain break for $l_{\mathrm{chain}}=4$ and $10$.
@@
As expected, for $\dot{\gamma}_{0}=0.1$ s-1 $<\dot{\gamma}_{\mathrm{sub.}}$ the chains never subdivide due to the magnetic dipole forces overpowering the weak viscous shear. At the other end of the strain rate range, $\dot{\gamma}_{0}=5$ s-1 $\gg\dot{\gamma}_{\mathrm{brk.}}$, for the majority of the cycle the viscous shear dominates over the magnetic effects and the particles act as mere tracers. It could be expected that when the instantaneous rate decreases the dipole interactions would regain relevance and a degree of particle aggregation would be expected. However, because the maximum applied rate is so large, the time interval where the instantaneous rate is sufficiently decreased is very small, and the particles have no time to aggregate, even at small deformations where they remain relatively close. The resulting chain-break conditions for $\dot{\gamma}_{0}=5$ s-1 are therefore non-representative of the magnetic-chain dynamics, but more of the relative movement of individual particles. On the other hand, if the maximum strain rate is not so large, such as $\dot{\gamma}_{0}=1$ s-1 $>\dot{\gamma}_{\mathrm{brk.}}$, the time interval where the dipole forces are drowned out is small, being insufficient to provoke complete chain breakdown but clearly sufficient for subdivision at large strains. On the other hand, for small $\gamma_{0}$, chain rotation is sufficient to comply with the imposed deformation and the chain remains unbroken. For $\dot{\gamma}_{0}=0.3$ and 0.7 s-1 the response is similar to $\dot{\gamma}=1$ s-1, except that, similarly to complete chain breakdown, the chain will only subdivide if a sufficient portion of the cycle is above the critical rate. If not, the chain does not have enough time to rotate past the critical angle, which is the case of the smaller chain ($l_{\mathrm{chain}}=4$) for $\dot{\gamma}_{0}=0.3$ s-1.
@@
In some other works23, 14, the beginning of viscous thinning may not agree with our critical strain, but this may be due to frequency dependence at larger $\omega$, larger particle concentrations that result in more complex microstructures or rheological differences of the continuous phase. The latter may be the case of the work of Wang et al., who report $v_{3}$ sign shifts occurring much earlier ($\gamma_{0}\sim 0.01$), but their continuous phase is deeply non-Newtonian, which should alter the microstructural dynamics as the viscous forces diverge significantly from the Newtonian case. In any case, this hypothesis should be corroborated with other similar samples before gathering definitive conclusions.
3.3.1 Magnetorheological alterations
@@
Again, there are no noticeable differences introduced by the carrier fluids. Moreover, the changes to the magnetic field, particle concentration and particle type do not alter the non-linear behaviour displayed by the Lissajous curves, but do, however, reduce the magnetorheological strength. Compared to the benchmark measurements, all alterations lead to a general decrease of both the storage and loss moduli. The least severe magnetorheological alteration is due to the field reduction to 50 mT, followed by the switch to the M270 particles and lastly, and most notoriously, the reduction of the particle concentration to 5 wt%. This weakening of the magnetorheological response leads to an approach of the experimental limits as the samples soften at high strain. Particularly, the 5 wtmeasurements display clear oscillations of the Lissajous curves, meaning the current setup is not suitable for the magnetorheological characterisation of weak MR fluids. In any case, these results agree with the steady shear data and the reasoning for the magnetorheological hierarchy between our measurements has already been discussed.
@@
Figure 11: Oscillatory shear data gathered with the NBa (left) and VBa (right) samples (see Table 1) at $\omega=0.1$ rad/s. (Top) raw Lissajous curves: blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also lightly plotted in the respective colours and the maximum measured stress is presented above each plot. (Bottom) First harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument inertia (corrected for the geometry’s gap-error).
@@
As mentioned above, the non-linear response evaluated at the tested strain values was not affected by the alterations. This is further corroborated by the FT-Chebyshev analysis. For the sake of conciseness, the data given by MITLaos is not shown here but is given as supplementary material (S3). The relevant moduli ($G^{\prime}_{M}$ and $G^{\prime}_{L}$) and dynamic viscosities ($\eta^{\prime}_{M}$ and $\eta^{\prime}_{L}$) all show the same tendency as the benchmark measurement (displayed in Figure 6), diminishing with strain (and strain rate) increase. $G^{\prime}_{M}$ and $G^{\prime}_{L}$ diverge past $\gamma_{0}>0.1$, with $G^{\prime}_{M}$ decreasing more acutely. For $\dot{\gamma}_{0}=0.1$ s-1, $\eta^{\prime}_{M}<\eta^{\prime}_{L}$ except for the 5wtmeasurement where $\eta^{\prime}_{M}\lesssim\eta^{\prime}_{L}$, which is probably due to low-torque errors. In any case, for all measurements, $\eta^{\prime}_{L}$ decreases faster than $\eta^{\prime}_{M}$. The third-order Chebyshev coefficients of the comparative measurements also display the same behaviour as the benchmark case. The elastic coefficient, $e_{3}$, always grows positive from zero at $\gamma_{0}=0.1$ (except for the 5 wtmeasurement, where $e_{3}/e_{1}\approx 5\%$, probably from experimental error) and the viscous counterpart, $v_{3}$, always crosses over from positive to negative between $10.1<\dot{\gamma}_{0}<$ 1, which may imply that the rheological and microstructural discussion is also valid for each of the alterations tested. Nonetheless, as we have seen, this sparse strain data set may hide important behaviours, and we believe the alterations tested should provoke some behavioural differences, as for example in the strain overshoot of $G^{\prime\prime}_{1}$.
@@ 4 Conclusions
Steady and oscillatory shear measurements were conducted with two blood analogues, one Newtonian and one viscoelastic, seeded with magnetic particles and under the influence of an external, uniform magnetic field perpendicular to the flow. The influence of magnetic field density, particle concentration and particle type were also evaluated by independently altering each parameter from a defined benchmark: 15wtof MyOne particles in a 250 mT magnetic field. The magnetic field was reduced to 50 mT, the particle concentration to 5 wtand lastly the MyOne participles were substituted with the M270.
@@
Under steady shear, the benchmark case returned a massive increase in low-shear viscosity and a clear shear-thinning behaviour, as the magnetically-induced microstructure progressively breaks down due to competing viscous shearing forces gaining relevance as the shear rate is increased. In oscillatory shear, the initial measurements displayed a practically frequency-independent response, but having $\omega\gtrsim 1$ rad/s gave rise to critical errors probably from inertial effects. A transition set was found with increasing strain: linear viscoelasticity to non-linear viscoelasticity to viscoplasticity, typical of field active fluids17, 14. The FT-Chebyshev analysis revealed the elastic softening was driven by the strain rate, whereas the viscous non-linearity was more complex. A finer amplitude sweep revealed a type III behaviour18, with a weak overshoot of $G_{1}^{\prime\prime}$ at the beginning of the non-linear regime that is associated with small structural rearrangements. The initial increase in $G_{1}^{\prime\prime}$ was accompanied by a positive contribution of the third-order viscous Chebyshev coefficient, $v_{3}>0$, which relates to viscous thickening driven by the strain rate. Increasing the strain, when $G_{1}^{\prime\prime}$ begins its descent, $v_{3}$ momentarily remained positive, which may be related to a shift of the driving mechanism of the viscous thinning non-linearity from the strain to the strain rate (at $\gamma_{0}\approx 0.5$).
@@
A magnetised-particle chain was modelled under oscillatory flow and a critical strain was found between $0.5\lesssim\gamma_{0}\lesssim 0.6$, above which chain rotation alone no longer complied with the imposed deformation and the chain was forced to subdivide. As such, the numerical results corroborated the experimental observation but more experimental data with similar magnetorheological fluids are necessary to confirm the hypothesis.
@@
All the alterations to the benchmark measurement resulted in a weaker magnetorheological response. In steady shear, the low-shear viscosity was reduced, particularly for the lesser particle concentration, and some evidence of apparent slip of the dispersed phase was noticed, meaning in future measurements both the geometry and the inferior plate must have significant surface roughness to avoid slip. In oscillatory shear, the alterations showed no significant differences in the non-linear behaviour, but finer sweeps are required to confirm this. The weakened magnetorheological response led to significant errors, which could compromise measurements with whole blood. The increase of the geometry radius seems to be the best solution to widen the experimental window but, for now, the geometrical restrictions associated with the commerciality available magnetorheological cells may be impeding a magnetorheological characterisation of blood under oscillatory shear.
@@ Acknowledgements
This work was financially supported by national funds through the FCT/MCTES (PIDDAC), under the project PTDC/EME-APL/3805/2021 (DOI 10.54499/PTDC/EME-APL/3805/2021), LA/P/0045/2020, UIDB/00532/2020 and UIDP/00532/2020, and the program Stimulus of Scientific Employment, Individual Support-2020.03203.CEECIND.
@@
S1 Apparent-slip with the CP20 MRD
Following the experimental campaign described at the end of Section 2, in Figure S1 are shown the flow curves, obtained with the CP20 MRD, of the magnetorheological benchmark measurements (15 wtof MyOne particles under a magnetic field of 250 mT) compared to the relevant parameter alterations (magnetic field density, particle concentration and particle type) and the unseeded blood analogues.
Figure S1: Flow curves obtained with the (left) NBa and (right) VBa, seeded with different particles (MyOne or M270), at different concentrations (5 and 15 wt%) and under magnetic fields of different densities (50 and 250 mT), with the CP20 MRD geometry at a commanded gap height of $h_{\mathrm{c}}=0.10$ mm. Experimental limits associated with low-torque issues and secondary flow onset are also presented.
The introduction of the magnetised particles returns a rheological response that seems practically independent of the applied field, the particle concentration and the particle type. In all cases there is an increase of the low-shear viscosity when compared to the unseeded analogues. Increasing the shear rate leads to a viscosity reduction, tending to that of the continuous phases’, but not resembling the usual, smooth shear-thinning behaviour. At middling shear rates there seems to be a tendency for a pseudo-Newtonian plateau that is more pronounced for the samples with the NBa as the continuous phase. This response at middling shear rates is reminiscent of previous reports of particle depletion near the smooth geometry walls. Because the continuous phase is less viscous, a substantial part of the imposed deformation is applied in this region, returning a lesser measured viscosity than if the particles were perfectly dispersed29, 30. One way to eliminate this apparent particle-slip is to employ geometries with enhanced surface roughness. We had at our disposal the serrated parallel-plate geometry for magnetorheological measurements, the PP20 MRD P2.
@@
Figure S2: Flow curves obtained with the (left) NBa and (right) VBa, without magnetic particles (0 wt%), with the PP20 MRD P2 geometry at different commanded gap heights ($0.05\leq h_{\mathrm{c}}\leq 0.30$ mm). Experimental limits associated with low-torque issues, maximum rotational velocity and secondary flows are also presented.
The flow curves clearly display a measured-viscosity dependence with the commanded gap height, yielding lesser viscosities the lower the gap. This can be explained either by slip effects or a gap-error42, but the latter seems more likely, probably arisen from a non-uniform serrated profile on the geometry’s surface. Having measured-viscosity data, $\eta_{\mathrm{m}}$, at different commanded gaps, $h_{\mathrm{c}}$, allows for an estimation of both the gap error, $\epsilon$, and the samples’ true viscosities, $\eta_{\mathrm{r}}$, by fitting the expression adapted from Kramer et al.:
@@
to the data at a specific shear rate. As can be seen, particularly for larger gaps, there are significant fluctuations in the flow curves. This means that there will also be variability in the gap-error and true viscosity estimation. Therefore, we chose to fit the expression to data gathered at multiple shear rates within a useful range ($10^{2}\leq\dot{\gamma}\leq 10^{3}$ s-1), given that at low and high shear the data is affected by experimental limitations, particularly large-gap data. The gap error was estimated as $\epsilon=123$ µm (standard error of 13 µm) and the flow curves obtained at a commanded gap of $h_{\mathrm{c}}=0.10$ mm are shown, corrected, in Figure S3. To compare the results of the PP20 MRD P2, in Figure S3 are also shown flow curves obtained with the CP20 MRD. The secondary flow and maximum rotational velocity limits are also gap-dependent and, because the gap-error is quite large, these were also corrected to the estimated true gap of the PP20 MRD P2, $h=h_{\mathrm{c}}+\varepsilon=223$ µm. The gap-error estimate seems to reasonably correct the PP20 MRD P2 results which are now similar to the data gathered with the CP20 MRD.
Figure S3: Flow curves obtained with the (left) NBa and (right) VBa, with no magnetic particles (0 wt%) nor field applied, with the PP20 MRD P2 ($h_{\mathrm{c}}=0.10$ mm) and CP20 MRD geometries. Experimental limits associated with low-torque issues, maximum rotational velocity and secondary flows are also presented (corrected for the gap error of the PP20 MRD P2).
Regarding the apparent-slip in magnetorheological measurements, in Figure S4 are shown the flow curves of the NBa benchmark sample (15 wtof MyOne particles under a magnetic field of 250 mT) gathered with the CP20 MRD and the PP20 MRD P2 (at a commanded gap of $h_{\mathrm{c}}=0.10$ mm and corrected for the gap-error). The PP20 MRD P2 flow curve shows a smooth shear-thinning response of the benchmark sample, with not evidence of the previously-noted pseudo-Newtonian plateau and with a much higher low-shear viscosity than with the CP20 MRD, almost one order of magnitude. Indeed, it seems that the serrated geometry was able to gather data free from the apparent-slip effect, returning the expected response. As such, we can assume the effects of apparent slip have been minimised; however, because our bottom plate is not detailed, nor was it treated with any surface finishing to enhance the roughness, we must beware, particularly because testing the gap variation seems to not be able to predict apparent slip in magnetorheological tests43.
Figure S4: Flow curves obtained with the NBa sample with 15 wtof MyOne particles, on the PP20 MRD P2 ($h_{\mathrm{c}}=0.10$ mm) and CP20 MRD geometries with an applied field density of $B=250$ mT. Experimental limits associated with low-torque issues, maximum rotational velocity and secondary flows are also presented (corrected for the gap error of the PP20 MRD P2).

### v2

@@ 3.3 Numerical analysis
Having determined the critical shear rates, we can define a strain rate range to test in oscillatory flow for either chain length. We aimed to test one strain rate for stable chains ($\dot{\gamma}_{0}<\dot{\gamma}_{\mathrm{sub.}}$), two within the intermediate range ($\dot{\gamma}_{\mathrm{sub.}}<\dot{\gamma}_{0}<\dot{\gamma}_{\mathrm{brk.}}$), and two past the breakdown threshold ($\dot{\gamma}_{0}>\dot{\gamma}_{\mathrm{brk.}}$ and $\dot{\gamma}_{0}\gg\dot{\gamma}_{\mathrm{brk.}}$). As such, for the chains with $l_{\mathrm{chain}}=4$ and 10, we defined, respectively: $\dot{\gamma}_{0}\in[0.1,0.3,0.7,1,5]$ s-1 and $\dot{\gamma}_{0}\in[0.01,0.1,0.7,1,5]$ s-1. Ten strain values were tested between $0.1\leq\gamma_{0}\leq 1$, running for 3 cycles ($t=3\times 2\pi/\omega$, sufficient to achieve a steady cycle) and the flow was driven by a top wall velocity $U=\dot{\gamma}_{0}\,h_{\mathrm{c}}\,\cos(\omega\,t)$, with $\omega=\dot{\gamma}_{0}/\gamma_{0}$.
@@
Having the particle positions, we want to probe for chain subdivision. In steady shear, we could simply analyse the interparticle distance as post chain-break the particles continuously diverge, but in oscillatory shear, this method alone is insufficient due to the time-dependent nature of the flow and because the minimum distance for chain break is somewhat arbitrary. Here we choose to consider a possible subdivision if, within the final simulated cycle, the distance between the central particle-pair exceeds 10% of their aggregated distance, $r/d_{\mathrm{p}}>1.1$, or if the critical angle is surpassed $\phi>33.5^{\circ}$. In Figure 11 are shown the passing conditions for chain break for $l_{\mathrm{chain}}=4$ and $10$.
@@
As expected, for $\dot{\gamma}_{0}=0.1$ s-1 $\dot{\gamma}_{\mathrm{sub.}}$ the chains never subdivide due to the magnetic dipole forces overpowering the weak viscous shear. At the other end of the strain rate range, $\dot{\gamma}_{0}=5$ s-1 $\gg\dot{\gamma}_{\mathrm{brk.}}$, for the majority of the cycle, the viscous shear dominates over the magnetic effects and the particles act as mere tracers. It could be expected that when the instantaneous rate decreases the dipole interactions would regain relevance and a degree of particle aggregation would be expected. However, because the maximum applied rate is so large, the time interval where the instantaneous rate is sufficiently decreased is very small, and the particles have no time to aggregate, even at small deformations where they remain relatively close. The resulting chain-break conditions for $\dot{\gamma}_{0}=5$ s-1 are therefore non-representative of the magnetic-chain dynamics, but more of the relative movement of individual particles. On the other hand, if the maximum strain rate is not so large, such as $\dot{\gamma}_{0}=1$ s-1 $\dot{\gamma}_{\mathrm{brk.}}$, the time interval where the dipole forces are drowned out is small, being insufficient to provoke complete chain breakdown but sufficient for subdivision at large strains. On the other hand, for small $\gamma_{0}$, chain rotation is enough to comply with the imposed deformation and the chain remains unbroken. For $\dot{\gamma}_{0}=0.3$ and 0.7 s-1 the response is similar to $\dot{\gamma}=1$ s-1, except that, similarly to complete chain breakdown, the chain will only subdivide if a sufficient portion of the cycle is above the critical rate. If not, the chain does not have enough time to rotate past the critical angle, which is the case of the smaller chain ($l_{\mathrm{chain}}=4$) for $\dot{\gamma}_{0}=0.3$ s-1.
@@
In some other workswang2021, li2004, the beginning of viscous thinning may not agree with our critical strain, but this may be due to frequency dependence at larger $\omega$, larger particle concentrations that result in more complex microstructures or rheological differences of the continuous phase. The latter may be the case of the work of wang2021, who report $v_{3}$ sign shifts occurring much earlier ($\gamma_{0}\sim 0.01$), but their continuous phase is deeply non-Newtonian, which should alter the microstructural dynamics as the viscous forces diverge significantly from the Newtonian case. In any case, this hypothesis should be corroborated with other similar samples before drawing definitive conclusions.
3.4 Magnetorheological conditions
@@
Again, no noticeable differences were introduced by the carrier fluids. Moreover, the changes to the magnetic field, particle concentration and particle type do not alter the non-linear behaviour displayed by the Lissajous curves, they do, however reduce the magnetorheological strength. Compared to the benchmark measurements, all alterations lead to a general decrease of both the storage and loss moduli. The least severe magnetorheological alteration is due to the field reduction to 50 mT, followed by the switch to the M270 particles, and most notoriously, the particle concentration reduction to 5 wt%. This weakening of the magnetorheological response leads to an approach of the experimental limits as the samples soften at high strain. Particularly, the 5 wt% measurements display clear oscillations of the Lissajous curves, meaning the current setup is not suitable for the magnetorheological characterisation of weak MR fluids. In any case, these results agree with the steady shear data, and the reasoning for the magnetorheological hierarchy between our measurements has already been discussed.
@@
Figure 12: Oscillatory shear data gathered with the NBa (left) and VBa (right) samples (see Table 1) at $\omega=0.1$ rad/s. (Top) raw Lissajous curves: blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also lightly plotted in the respective colours and the maximum measured stress is presented above each plot. (Bottom) First-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument inertia (corrected for the geometry’s gap-error).
@@
As mentioned above, the alterations did not affect the non-linear response evaluated at the tested strain values. This is further corroborated by the FT-Chebyshev analysis. For the sake of conciseness, the data given by MITLaos is not shown here but is given as supplementary material (S6). The relevant moduli ($G^{\prime}_{M}$ and $G^{\prime}_{L}$) and dynamic viscosities ($\eta^{\prime}_{M}$ and $\eta^{\prime}_{L}$) all show the same tendency as the benchmark measurement (displayed in Figure 6), diminishing with strain (and strain rate) increase. $G^{\prime}_{M}$ and $G^{\prime}_{L}$ diverge past $\gamma_{0}>0.1$, with $G^{\prime}_{M}$ decreasing more acutely. For $\dot{\gamma}_{0}=0.1$ s-1, $\eta^{\prime}_{M}<\eta^{\prime}_{L}$ except for the 5wt% measurement where $\eta^{\prime}_{M}\lesssim\eta^{\prime}_{L}$, which is probably due to low-torque errors. In any case, for all measurements, $\eta^{\prime}_{L}$ decreases faster than $\eta^{\prime}_{M}$. The third-order Chebyshev coefficients of the comparative measurements also display the same behaviour as the benchmark case. The elastic coefficient, $e_{3}$, always grows positive from zero at $\gamma_{0}=0.1$ (except for the 5 wt% measurement, where $e_{3}/e_{1}\approx 5\%$, probably from experimental error). The viscous counterpart, $v_{3}$, always crosses over from positive to negative between $0.1<\dot{\gamma}_{0}<1$, which may imply that the rheological and microstructural discussion is also valid for each of the alterations tested. Nonetheless, as we have seen, this sparse strain data set may hide important behaviours, and we believe the alterations tested should provoke some behavioural differences, as for example in the strain overshoot of $G^{\prime\prime}_{1}$.
@@ 4 Conclusions
Steady and oscillatory shear measurements were conducted using two blood analogues, one Newtonian and one viscoelastic, both seeded with magnetic particles and subjected to an external, uniform magnetic field perpendicular to the flow. The influence of magnetic field density, particle concentration and particle type was also evaluated by independently altering each parameter from a defined benchmark: 15wt% of MyOne particles in a 250 mT magnetic field. The magnetic field was reduced to 50 mT, the particle concentration to 5 wt%, and the MyOne particles were replaced with the M270.
@@
Under steady shear, the benchmark case returned a massive increase in low-shear viscosity and a clear shear-thinning behaviour, as the magnetically-induced microstructure progressively breaks down due to competing viscous shearing forces gaining relevance as the shear rate increases. The response is well described by the Casson model, even though the yield stress could not be directly observable in the flow curves, possibly due to apparent slip on the bottom plate of the rheometer. In oscillatory shear, the initial measurements displayed a practically frequency-independent response; however, having $\omega\gtrsim 1$ rad/s gave rise to critical errors likely due to inertial effects. A transition set was observed with increasing strain: linear viscoelasticity to non-linear viscoelasticity to viscoplasticity, typical of field-active fluidsparthasarathy1999, li2004. The FT-Chebyshev analysis revealed that the elastic softening was driven by the strain rate, whereas the viscous non-linearity was more complex. A finer amplitude sweep revealed a type III behaviourhyun2002, with a weak overshoot of $G_{1}^{\prime\prime}$ at the beginning of the non-linear regime associated with small structural rearrangements. The initial increase in $G_{1}^{\prime\prime}$ was accompanied by a positive contribution of the third-order viscous Chebyshev coefficient, $v_{3}>0$, which relates to viscous thickening driven by the strain rate. Increasing the strain, when $G_{1}^{\prime\prime}$ begins its descent, $v_{3}$ momentarily remained positive, which may be related to a shift of the driving mechanism of the viscous thinning non-linearity from the strain to the strain rate (at $\gamma_{0}\approx 0.5$).
@@
A magnetised-particle chain was modelled under oscillatory flow, and a critical strain was found between $0.5\lesssim\gamma_{0}\lesssim 0.6$, above which chain rotation alone no longer complied with the imposed deformation and the chain was forced to subdivide. As such, the numerical results corroborate the experimental observation, but more experimental data with similar magnetorheological fluids are necessary to confirm the hypothesis.
@@
All the alterations to the benchmark measurement resulted in a weaker magnetorheological response. In steady shear, the low-shear viscosity was reduced, particularly for the lesser particle concentration, and some evidence of apparent slip of the dispersed phase was noticed, meaning in future measurements both the geometry and the bottom plate must have significant surface roughness to avoid slip. In any case, the experimental data generally collapsed onto a single master curve dependent on the Mason number. In oscillatory shear, the alterations showed no significant differences in the non-linear behaviour, but finer sweeps are required to confirm this. The weakened magnetorheological response led to significant errors, which could compromise measurements with whole blood. The increase in the geometry radius is the best solution to widen the experimental window, but, for now, the geometrical restrictions associated with commercially available magnetorheological cells may be impeding a magnetorheological characterisation of blood under oscillatory shear.
@@ Acknowledgements
This work was financially supported by national funds through the FCT/MCTES (PIDDAC), under the project PTDC/EME-APL/3805/2021 (DOI 10.54499/PTDC/EME-APL/3805/2021), by the FEDER funds through COMPETE2030 and with financial support of FCT, I.P., within the framework of the project PROMisH - Potential of Rheologically-Optimized Magnetic Emulsions in Hemotherapy, with nr. 16207, operation code at the Funds Platform COMPETE2030-FEDER-00736400 and FCT code 2023.17217.ICDT With DOI https://doi.org/10.54499/2023.17217.ICDT., LA/P/0045/2020, UIDB/00532/2020 and UIDP/00532/2020, and the program Stimulus of Scientific Employment, Individual Support-2020.03203.CEECIND. R.R. acknowledges FCT for funding support under scholarship No. 2025.03368.BD.
@@
S1 Apparent slip with the CP20 MRD
Following the experimental campaign described at the end of Section 2, in Figure S1 are shown the flow curves, obtained with the CP20 MRD, of the magnetorheological benchmark measurements (15 wt% of MyOne particles under a magnetic field of 250 mT) compared to the relevant parameter alterations (magnetic field density, particle concentration and particle type) and the unseeded blood analogues.
Figure S1: Viscosity curves obtained with the NBa (left column) and VBa (right column), seeded with MyOne particles at 15 wt% under the influence of 50 and 250 mT (top row), with MyOne particles at 5 and 15 wt% under 250 mT (middle row), and with MyOne and M270 particles at 15 wt% under 250 mT (bottom row). Adjusted low-torque and secondary flow limits are also shown in each graph. Data gathered with the CP20 MRD geometry.
The introduction of the magnetised particles returns a rheological response that appears practically independent of the applied field, particle concentration, and particle type. In all cases, there is an increase in the low-shear viscosity when compared to the unseeded analogues. Increasing the shear rate leads to a viscosity reduction, tending to that of the continuous phases, but not resembling the usual, smooth shear-thinning behaviour. At middling shear rates, there seems to be a tendency for a pseudo-Newtonian plateau that is more pronounced for the samples with the NBa as the continuous phase. This response at middling shear rates is reminiscent of previous reports of particle depletion near the smooth geometry walls. Because the continuous phase is less viscous, a substantial part of the imposed deformation is applied in this region, returning a lesser measured viscosity than if the particles were perfectly dispersedbarnes1995, vicente2004. One way to eliminate this apparent particle slip is to employ geometries with enhanced surface roughness, as the serrated parallel-plate geometry for magnetorheological measurements, the PP20 MRD P2.
@@
Unseeded
Figure S2: Flow curves obtained with the (left) NBa and (right) VBa, without magnetic particles (0 wt%), with the PP20 MRD P2 geometry at different commanded gap heights ($0.05\leq h_{\mathrm{c}}\leq 0.30$ $m$). Experimental limits associated with low-torque issues, maximum rotational velocity and secondary flows are also presented.
The flow curves clearly display a measured-viscosity dependence on the commanded gap height, yielding lower viscosities as the gap decreases. This can be explained either by slip effects or a gap-errorvleminckx2016, but the latter seems more likely as detailed or roughened geometries allow the sample to flow within the solid structurescarotenuto2013, nickerson2005. Having measured-viscosity data, $\eta_{\mathrm{m}}$, at different commanded gaps, $h_{\mathrm{c}}$, allows for an estimation of both the gap error, $\epsilon$, and the samples’ true viscosities, $\eta_{\mathrm{r}}$, by fitting the expression adapted from kramer1987:
@@
to the data at a specific shear rate. As can be seen, particularly for larger gaps, there are significant fluctuations in the flow curves. This means that there will also be variability in the gap-error and true viscosity estimation. Therefore, we chose to fit the expression to data gathered at multiple shear rates within a useful range ($10^{2}\leq\dot{\gamma}\leq 10^{3}$ s-1), given that at low and high shear the data is affected by experimental limitations, particularly large-gap datarodrigues2025. The gap error was estimated as $\epsilon=123$ µm (standard error of 13 µm) and the flow curves obtained at a commanded gap of $h_{\mathrm{c}}=0.10$ mm are shown, corrected, in Figure S3. To compare the results of the PP20 MRD P2, in Figure S3 flow curves obtained with the CP20 MRD are also shown. The secondary flow and maximum rotational velocity limits are also gap-dependent and, because the gap-error is quite large, these were also corrected to the estimated true gap of the PP20 MRD P2, $h=h_{\mathrm{c}}+\varepsilon=223$ µm. The gap-error estimate seems to reasonably correct the PP20 MRD P2 results, which are now similar to the data gathered with the CP20 MRD.
Unseeded
Figure S3: Flow curves obtained with the (left) NBa and (right) VBa, with no magnetic particles (0 wt%) nor field applied, with the PP20 MRD P2 ($h_{\mathrm{c}}=0.10$ $m$) and CP20 MRD geometries. Experimental limits associated with low-torque issues, maximum rotational velocity and secondary flows are also presented (corrected for the gap error of the PP20 MRD P2).
Regarding the apparent slip in magnetorheological measurements, in Figure S4 are shown the flow curves of the NBa benchmark sample (15 wt% of MyOne particles under a magnetic field of 250 mT) gathered with the CP20 MRD and the PP20 MRD P2 (at a commanded gap of $h_{\mathrm{c}}=0.10$ mm and corrected for the gap-error). The PP20 MRD P2 flow curve shows a smooth shear-thinning response of the benchmark sample, with no evidence of the previously-noted pseudo-Newtonian plateau and with a much higher low-shear viscosity than with the CP20 MRD (almost one order of magnitude). Indeed, it seems that the serrated geometry was able to gather data free from the apparent slip effect, returning the expected response. As such, we can assume the effects of apparent slip have been minimised; however, because our bottom plate is not detailed, nor was it treated with any surface finishing to enhance the roughness, we must beware, particularly because testing the gap variation seems not to be able to predict apparent slip in magnetorheological testsjonkkari2012.
15_MyOne_250 (NBa)
Figure S4: Flow curves obtained with the NBa sample with 15 wt% of MyOne particles, on the PP20 MRD P2 ($h_{\mathrm{c}}=0.10$ $m$) and CP20 MRD geometries with an applied field density of $B=250$ mT. Experimental limits associated with low-torque issues, maximum rotational velocity and secondary flows are also presented (corrected for the gap error of the PP20 MRD P2).